# Learning Chemical Reaction Representation with Reactant-Product Alignment

### v1

@@ 4.2. Incorporating Reaction Conditions
Drawing inspiration from multimodal conditional image generation works such as T2IAdapter (Mou et al., 2024) or ControlNet (Zhang et al., 2023), which adeptly integrate multimodal information and effectively leverage prior research to enhance model performance, we have chosen to implement an adapter structure for incorporating reaction conditions. This approach allows us to seamlessly assimilate these conditions without necessitating modifications to the underlying architecture of the Atom Aligned Encoder. Let us assume that the reaction condition for a given reaction has been encoded into a feature matrix , where denotes the number of features and signifies the dimension of each feature. For each block within the Atom Aligned Encoder, we utilize multi-head attention to integrate the reaction condition information into its output node features. Subsequently, we employ these node features, now imbued with reaction condition information, to generate the output edge features. Mathematically, with the incorporation of chemical reaction conditions, the output node features and edge features of the -th block in Atom Aligned Encoder is modified as follows:
@@
4.3. Reaction-Center-Aware Decoders
The decoder takes the encoded node features and from Atom-Aligned Encoder with blocks as input and generates outputs in various formats depending on the task at hand. In this section, we introduce decoder architectures tailored for both sequential generation tasks and tasks with a single output. Given that reaction centers record the key functional groups involved in chemical reactions and play a decisive role in their properties (Keto et al., 2024), we have designed a Reaction-Center-Aware Decoder that explicitly integrates information about reaction centers into the chemical reaction representations applied to downstream tasks.
We first introduce the Reaction-Center-Aware (RC-aware) cross-attention mechanism, which is adapted from the local-global decoder of the Retroformer (Wan et al., 2022). The RC-aware cross-attention is a specialized attention mechanism where half of the attention heads function identically to the standard attention, while the other half is restricted to accessing only the node features of the reaction centers . The formulation for attention heads within the RC-aware cross-attention mechanism, which are constrained to accessing only the reaction centers, is articulated as follows:
@@ 4.3. Reaction-Center-Aware Decoders
where is learnable parameter, represents the concatenation of features and is the number of heads.
Then for sequential generation tasks, we replace the cross-attention layers of vanilla transformer decoder (Vaswani et al., 2017) with the RC-aware cross-attention, thereby customizing our decoder. In the case of tasks that require a numerical output, such as chemical reaction yield prediction, we utilize a learnable query vector within the RC-aware cross-attention mechanism to derive a reaction-level representation. Subsequently, this representation is fed into a feed-forward network to generate the final output.
5. Experiments
@@
| Parrot-LM_E (Wang et al., 2023) | 92.12 | 94.91 | 95.97 | 97.28 | 44.20 | 59.23 | 63.71 | 66.16 | 46.47 | 62.04 | 68.19 | 73.93 | ||
| GCNN (Maser et al., 2021) | 90.59 | 91.80 | 92.40 | 93.39 | 32.39 | 46.02 | 52.51 | 61.06 | 35.84 | 46.37 | 50.61 | 55.99 | ||
| Reagent Transformer (Andronov et al., 2023) | 89.80 | 93.30 | 94.52 | 95.88 | 37.97 | 51.29 | 57.47 | 65.44 | 39.11 | 55.20 | 61.83 | 69.80 | ||
| FPRCR (Gao et al., 2018) | 91.22 | 92.82 | 93.57 | 94.74 | 38.51 | 49.72 | 54.64 | 61.56 | 37.54 | 47.95 | 52.69 | 58.48 | ||
@@
| GCNN (Maser et al., 2021) | 12.81 | 21.95 | 26.40 | 32.17 |
| FPRCR (Gao et al., 2018) | 16.90 | 26.38 | 31.16 | 36.96 |
| Reagent Transformer (Andronov et al., 2023) | 22.74 | 34.09 | 39.36 | 46.01 |
| Parrot-LM-E (Wang et al., 2023) | 27.42 | 41.86 | 46.98 | 50.95 |
@@
| T5Chem-pretrained (Lu and Zhang, 2022) | 26.2 | 39.4 | 45.0 | 51.6 |
| GCNN (Maser et al., 2021) | 4.92 | 12.66 | 18.35 | 25.19 |
| Reagent Transformer (Andronov et al., 2023) | 19.20 | 27.24 | 30.81 | 35.20 |
| T5Chem-from-scratch (Lu and Zhang, 2022) | 17.50 | 28.00 | 32.8 | 38.70 |
| Ours | 26.84 | 39.58 | 44.61 | 50.25 |
@@
To demonstrate the capability of our model to extract potent reaction embeddings and apply them to a variety of downstream tasks, we conducted extensive experiments, including predictions of reaction condition combinations, reaction yield, and reaction selectivity. For ease of understanding, we have depicted the pipelines for different tasks in Fig 2.
5.1. Reaction Condition Combination Prediction/Generation
@@
Dataset. We use USPTO_CONDITION dataset to evaluate performance of our model on reaction condition combination prediction task. This dataset comprises 680,741 reactions, with each reaction condition being consistently composed of one catalyst, two reagents, and two solvents. We have directly utilized the pre-processed data from Wang et al. and further employed RXNMapper (Schwaller et al., 2021a) to augment the chemical reactions with atom mapping. For reaction condition generation tasks, we use USPTO_500MT dataset for evaluation. We generate and sort the reagents according to a certain rule, whose detail is displayed in Appendix C.2. The tokenization on the SMILES representations of the reagents is aligned with the work (Schwaller et al., 2019).
Metric. In this section, we assess the predictive performance for the whole reaction condition combinations as well as for each constituent element within the reaction conditions. We use the conventional top- accuracy to evaluate the performance of model. A prediction is considered as correct if and only if all the molecules of it are correctly predicted. The order of the constituents is not one of the criteria for determining whether a prediction is correct. When a reaction in the test set has multiple recorded reaction combinations, a prediction is considered correct if it is completely consistent with any one of them.
Baselines. We compare our method against four baselines for reaction condition prediction tasks. GCNN (Maser et al., 2021) uses massage-passing networks to extract reaction representation and then predicts the reaction condition combinations. FPRCR (Gao et al., 2018) utilizes the fingerprints of reactants, products, and the components of the reaction conditions that have already been predicted as inputs to predict the next composition of reaction condition combinations. Parrot-LM-E (Wang et al., 2023) and Reagent Transformer (Andronov et al., 2023) are two transformer-based models that have been developed, respectively, based on the checkpoint of BERT (Devlin et al., 2019) for Parrot-LM-E and the checkpoint of the Molecular Transformer (Schwaller et al., 2019) for Reagent Transformer. Note that the Parrot-LM-E and FPRCR are specifically designed for reaction condition combinations with a fixed number of components, so they have not been applied to the USPTO_500MT dataset. We also report the experimental results of T5chem (Lu and Zhang, 2022), a text-to-text transformer model, on the USPTO_500MT dataset. The results from a model that was finetuned from a checkpoint pretrained on a large-scale reaction dataset, as well as the results from a model trained from scratch, have both been reported by us.
@@
Performance Evaluation. The USPTO_CONDITION dataset’s result is summarized in Table 3 and Table 1. And the results of USPTO_500MT is summarized in Table 3. From the table we can find that on USPTO_CONDITION dataset, our model achieves a top-1 accuracy of of 34.30%, a top-5 accuracy of 52.82% and a top-10 accuracy of 59.22%, surpassing the strongest baseline Parrot-LM_E by 6.88%, 5.84% and 8.27% respectively. When evaluating the prediction accuracy for each type of component within reaction conditions, it is observed that our model significantly outperforms all the baseline models across all metrics. Particularly in the prediction of solvents, our model attains a top-1 accuracy of 50.45% and a top-10 accuracy of 79.18%, surpassing the strongest baseline by 6.25% and 13.02%, respectively. On USPTO_500MT dataset, our model achieves a top-1 overall accuracy of 26.84% and an top-10 overall accuracy of 50.25%, which exceeds the strongest baseline that did not utilize pretraining by 7.64% and 11.55%. Additionally, it can be observed that our model exhibits performance almost equivalent to that of the T5Chem model trained on a large-scale reaction dataset. The aforementioned performance indicates that our model is capable of extracting robust reaction representations for both predictive and generative tasks.
@@
| DRFP (Probst et al., 2022) | 7.9492 | 11.3285 | 0.8273 | 9.0878 | 13.5990 | 0.7480 | 10.0901 | 15.8577 | 0.6814 | 12.7572 | 19.1300 | 0.4769 | |||
| Chemprop (Heid et al., 2024) | 8.5883 | 12.3130 | 0.7960 | 10.5984 | 14.2913 | 0.7217 | 10.4930 | 15.4690 | 0.6969 | 14.5839 | 20.4564 | 0.4018 | |||
| YieldBert (Schwaller et al., 2021c) | 7.5416 | 11.2156 | 0.8308 | 7.4349 | 10.8098 | 0.8408 | 9.6488 | 15.2584 | 0.7051 | 13.5600 | 19.3862 | 0.4627 | |||
| T5Chem (Lu and Zhang, 2022) | 7.1283 | 11.2242 | 0.8385 | 6.7693 | 10.4041 | 0.8801 | 9.0982 | 14.3431 | 0.7665 | 13.4069 | 19.7521 | 0.6051 | |||
@@
| DRFP (Probst et al., 2022) | 4.0995 0.1191 | 6.2424 0.2636 | 0.9474 0.0050 |
| Chemprop (Heid et al., 2024) | 4.6430 0.1405 | 6.4306 0.1938 | 0.9441 0.0042 |
| YieldBert (Schwaller et al., 2021c) | 3.5532 0.1600 | 5.4480 0.3240 | 0.9599 0.0050 |
| T5Chem (Lu and Zhang, 2022) | 3.5059 0.1562 | 5.3181 0.2482 | 0.9662 0.0034 |
@@
5.2. Reaction Yield Prediction
Dataset. We use Buchwald-Hartwig dataset (Ahneman et al., 2018) to evaluate the performance of our model on reaction yield prediction task. The dataset provides 10 random splits and four ligand-based out-of-sample splits for evaluation. The test sets under the four out-of-sample data split contain reaction additives which are not included in the train sets. We use the raw data and data split provided by Probst et al. and add the atom mapping for reactions according to the reaction template. In this study, we have standardized the yields to a range of 0 to 100. The statistical information of different splits is summarized in Appendix A.
Baselines. We use four powerful baselines for comparison. The first is DRFP (Probst et al., 2022), a kind of chemical reaction fingerprint for yield prediction. The second is Chemprop (Heid et al., 2024), a kind of Message Passing Network encoding multi-molecules or reactions. The third is YieldBert (Schwaller et al., 2021c), a transformer pretrained on Pistachio dataset (Mayfield et al., 2017) using self-supervised tasks. And the fourth is T5Chem (Lu and Zhang, 2022), a language model pretrained on PubMed dataset (Kim et al., 2020) with a self-supervised task and USPTO_500MT dataset (Lu and Zhang, 2022) with five different supervised tasks including reaction yield prediction.
Implementation Details. It should be noted that the number of distinct reagents in the reaction conditions of Buchwald-Hartwig dataset is less than 50, which makes it challanging to train a reaction condition encoder from scratch. Thus we use a lightweight pretrained molecular encoder (Hu et al., 2020b) as our reaction condition encoder. For detailed information, please refer to Appendix B.4.
@@ 5.2. Reaction Yield Prediction
| MFF (Sandfort et al., 2020) | 2.2923 0.0400 | 2.9388 0.0393 | 0.5891 0.0095 |
| DRFP (Probst et al., 2022) | 0.9435 0.0293 | 1.2943 0.0387 | 0.9203 0.0042 |
| Chemprop (Heid et al., 2024) | 0.3593 0.0099 | 0.5396 0.0299 | 0.9861 0.0015 |
| RXNFP (Schwaller et al., 2021b) | 0.3744 0.0094 | 0.5378 0.0204 | 0.9862 0.0010 |
| T5Chem (Lu and Zhang, 2022) | 0.6272 0.0184 | 0.8213 0.0226 | 0.9822 0.0011 |
@@
| MFF (Sandfort et al., 2020) | 0.1421 0.0093 | 0.2122 0.0169 | 0.9055 0.0119 |
| DRFP (Probst et al., 2022) | 0.1481 0.0094 | 0.2120 0.0147 | 0.9056 0.0117 |
| Chemprop (Heid et al., 2024) | 0.1626 0.0118 | 0.2283 0.0174 | 0.8907 0.0130 |
| RXNFP (Schwaller et al., 2021b) | 0.1650 0.0111 | 0.2313 0.0197 | 0.8872 0.0193 |
| T5Chem (Lu and Zhang, 2022) | 0.1662 0.0108 | 0.2417 0.0182 | 0.8920 0.0121 |
@@
| Full Version | 26.84 | 39.58 | 44.61 | 50.25 | 3.6331 0.1259 | 5.5649 0.2839 | 0.9581 0.0049 | |
| - Atom Aligned Encoder | 26.01 | 38.79 | 43.99 | 50.19 | 3.7289 0.1291 | 5.7318 0.2520 | 0.9556 0.0044 | |
| - Reaction-Center-Aware Decoders | 26.29 | 39.14 | 43.91 | 49.82 | 3.6967 0.1710 | 5.7845 0.3948 | 0.9547 0.0066 | |
Performance Evaluation. The average results of ten random splits of Buchwald-Hartwig dataset is summarized in Table 5 and the results of four out-of-sample splits are summarized in Table 4. Table 5 indicates that our model has achieved an average of 0.958 and an average MAE of 3.633 under ten random splits, placing it in the third position among all the compared methods. However, it is important to note that our model, which has not been pretrained on large-scale reaction data, is not significantly outperformed by two deep learning models that have. The difference in between our model and the strongest baseline is less than 0.01, and the gap in terms of MAE is less than 0.1. In the out-of-sample settings, our model demonstrates superior performance under various evaluation metrics. For instance, it attains an score of 0.898, surpassing the strongest baseline by 0.060. Notably, our model outperforms other baselines in all metrics, excluding the metric in the Test4 data split, by a significant margin. Furthermore, our model records an approximate 2.000 improvement in RMSE across all data splits.
@@
5.3. Reaction Selectivity Prediction
Datasets. We use the C-H functionalization dataset (Li et al., 2020) created by Li et al. to demonstrate the regio-selectivity prediction performance, and use the experimental dataset (Zahrt et al., 2019; Li et al., 2023) regarding chiral phosphoric acid–catalyzed thiol addition to N-acylimines created by Zahrt et al. to illustrate the enantioselectivity predictive performance. There are 6114 chemical reactions in C-H functionalization Dataset, and thiol addition Dataset contains 43 catalysts and 5 5 reactants combination, which form 1075 reactions. These datasets was randomly divided into a training set and a test set with a ratio of 7:3 for 10 times.
Baseline. Considering that the reaction selectivity prediction task is fundamentally a regression problem, we have adapted three deep-learning models originally designed for reaction yield prediction, Chemprop (Heid et al., 2024), T5chem (Lu and Zhang, 2022) and RXNFP (Schwaller et al., 2021b), to this task. Additionally, two fingerprint-based methods are compared as well. In detail, DRFP (Probst et al., 2022) calculate the fingerprint of the symmetric difference between the n-grams of reactants and products, while MFF (Sandfort et al., 2020) leverages multiple fingerprint features of all the molecules in one reaction as input of the regressor.
Performance Evaluation. The results on two datasets are summarized in Table 6 and Table 7. On C-H functionalization Dataset, our model achieves the best performance among all the compared methods across all metrics, with an average MAE of 0.327, an average RMSE of 0.511, and an average of 0.988 Notably, our models outperforms two baselines RXNFP and T5chem, which are pretrained on large-scale reaction dataset. These results demonstrate that our model exhibits architectural superiority and a stronger understanding of chemical reactions compared to existing methods.

### v2

@@ 4.2 Incorporating Reaction Conditions
Drawing inspiration from multimodal conditional image generation works such as T2IAdapter (Mou et al., 2024) or ControlNet (Zhang et al., 2023), which adeptly integrate multimodal information and effectively leverage prior research to enhance model performance, we have chosen to implement an adapter structure for incorporating reaction conditions. This approach allows us to seamlessly assimilate these conditions without necessitating modifications to the underlying architecture of the Atom Aligned Encoder. Let us assume that the reaction condition for a given reaction has been encoded into a feature matrix , where denotes the number of features and signifies the dimension of each feature. For each block within the Atom Aligned Encoder, we utilize multi-head attention to integrate the reaction condition information into its output node features. Subsequently, we employ these node features, now imbued with reaction condition information, to generate the output edge features. Mathematically, with the incorporation of chemical reaction conditions, the output node features and edge features of the -th block in Atom Aligned Encoder are modified as follows:
@@
4.3 Reaction-Center-Aware Decoders
The decoder takes the encoded node features and from Atom-Aligned Encoder with blocks as input and generates outputs in various formats depending on the task at hand. In this section, we introduce decoder architectures tailored for both sequential generation tasks and tasks with a single output. Given that reaction centers record key functional groups involved in chemical reactions and play a decisive role in their properties (Keto et al., 2024), we have designed a Reaction-Center-Aware Decoder that explicitly integrates information about reaction centers into chemical reaction representations applied to downstream tasks.
We first introduce the Reaction-Center-Aware (RC-aware) cross-attention mechanism, which is adapted from the local-global decoder of the Retroformer (Wan et al., 2022). RC-aware cross-attention is a specialized attention mechanism in which half of the attention heads function identically to standard attention, while the other half is restricted to accessing only the node features of the reaction centers . The formulation for attention heads within the RC-aware cross-attention mechanism, which are constrained to accessing only the reaction centers, is articulated as follows:
@@ 4.3 Reaction-Center-Aware Decoders
where is a learnable parameter, represents the concatenation of features and is the number of heads.
Then for sequential generation tasks, we replace the cross-attention layers of the vanilla transformer decoder (Vaswani et al., 2017) with RC-aware cross-attention, thus customizing our decoder. In the case of tasks that require a numerical output, such as chemical reaction yield prediction, we utilize a learnable query vector within the RC-aware cross-attention mechanism to derive a reaction-level representation. Subsequently, this representation is fed into a feed-forward network to generate the final output.
5 Experiments
@@
| Parrot-LM_E | 92.12 | 94.91 | 95.97 | 97.28 | 44.20 | 59.23 | 63.71 | 66.16 | 46.47 | 62.04 | 68.19 | 73.93 | ||
| GCNN | 90.59 | 91.80 | 92.40 | 93.39 | 32.39 | 46.02 | 52.51 | 61.06 | 35.84 | 46.37 | 50.61 | 55.99 | ||
| Reagent Transformer | 89.80 | 93.30 | 94.52 | 95.88 | 37.97 | 51.29 | 57.47 | 65.44 | 39.11 | 55.20 | 61.83 | 69.80 | ||
| FPRCR | 91.22 | 92.82 | 93.57 | 94.74 | 38.51 | 49.72 | 54.64 | 61.56 | 37.54 | 47.95 | 52.69 | 58.48 | ||
| T5Chem | 88.60 | 92.38 | 93.68 | 95.07 | 30.30 | 46.66 | 54.63 | 64.57 | 32.71 | 47.98 | 54.86 | 63.25 | ||
@@
| GCNN | 12.81 | 21.95 | 26.40 | 32.17 |
| FPRCR | 16.90 | 26.38 | 31.16 | 36.96 |
| Reagent Transformer | 22.74 | 34.09 | 39.36 | 46.01 |
| Parrot-LM-E | 27.42 | 41.86 | 46.98 | 50.95 |
| T5Chem | 16.42 | 27.75 | 33.53 | 40.79 |
@@
| T5Chem-pretrained | 25.36 | 38.12 | 43.42 | 49.54 |
| GCNN | 4.74 | 12.36 | 17.94 | 24.58 |
| Reagent Transformer | 17.88 | 25.47 | 28.78 | 33.27 |
| T5Chem-from-scratch | 17.32 | 28.21 | 32.98 | 38.76 |
| Ours | 26.91 | 39.80 | 44.82 | 50.73 |
@@
To demonstrate the capability of our model to extract potent reaction embeddings and its adaptability to various downstream tasks, we conducted extensive experiments, including predictions of reaction condition combinations, reaction yield, and reaction selectivity. For ease of understanding, we have depicted the pipelines for different tasks in Fig 9.
5.1 Reaction Condition Combination Prediction/Generation
@@
Dataset. We use the USPTO_CONDITION dataset to evaluate the performance of our model on the reaction condition combination prediction task. This dataset comprises 680,741 reactions, with each reaction condition consistently consisting of one catalyst, two reagents, and two solvents. We have directly used the data processed by Wang et al. and further employed RXNMapper (Schwaller et al., 2021a) to augment chemical reactions with atom mapping. For the reaction condition combination generation task, we use the USPTO_500MT dataset for evaluation. We generate and sort the reagents according to a certain rule, whose details are displayed in Appendix D.2. The tokenization on the SMILES representations of the reagents is aligned with the work (Schwaller et al., 2019). We use the top- accuracy to evaluate both prediction and generation tasks; the details are illustrated in Appendix B.2.
Baselines. We choose T5Chem (Lu & Zhang, 2022), GCNN (Maser et al., 2021) and Reagent Transformer (Andronov et al., 2023) as our baselines on both the USPTO_500MT and the USPTO_CONDITION datasets. For the USPTO_CONDITION dataset, we compared two more baselines, Parrot-LM_E (Wang et al., 2023) and FPRCR (Gao et al., 2018), which are designed for reaction condition combinations with a fixed number of components. Additionally, we reported the performance of T5Chem finetuned from the pretrained checkpoint on the USPTO_500MT dataset. The details of the baselines are in Appendix E.
@@
Performance Evaluation. The results of the USPTO_CONDITION dataset are summarized in Table 3 and Table 1. And the results of USPTO_500MT are summarized in Table 3. From the table, we can find that in the USPTO_CONDITION dataset, our model achieves a top-1 accuracy of 34.30%, a top-5 accuracy of 52.82%, and a top-10 accuracy of 59.22%, surpassing the strongest baseline Parrot-LM_E by 6.88%, 5.84%, and 8.27%, respectively. When evaluating the prediction accuracy for each type of component within reaction conditions, it is observed that our model significantly outperforms all baseline models across all metrics. Particularly in the prediction of solvents, our model attains a top-1 accuracy of 50.45% and a top-10 accuracy of 79.18%, surpassing the strongest baseline by 6.25% and 13.02%, respectively. On the USPTO_500MT dataset, our model achieves a top-1 overall accuracy of 26.91% and a top-10 overall accuracy of 50.73%, which exceeds the strongest baseline that did not utilize pretraining by 9.03% and 11.93%. It can be observed that our model outperforms the T5Chem model pretrained on a large-scale reaction dataset. The aforementioned performance indicates that our model is capable of extracting robust reaction representations for both predictive and generative tasks.
@@
| MFF | 9.7774 | 14.2630 | 0.7263 | 9.7732 | 14.2123 | 0.7248 | 9.6202 | 14.9177 | 0.7181 | 14.7023 | 19.8501 | 0.4367 | |||
| DRFP | 7.9492 | 11.3285 | 0.8273 | 9.0878 | 13.5990 | 0.7480 | 10.0901 | 15.8577 | 0.6814 | 12.7572 | 19.1300 | 0.4769 | |||
| Chemprop | 8.5883 | 12.3130 | 0.7960 | 10.5984 | 14.2913 | 0.7217 | 10.4930 | 15.4690 | 0.6969 | 14.5839 | 20.4564 | 0.4018 | |||
| YieldBert | 7.5416 | 11.2156 | 0.8308 | 7.4349 | 10.8098 | 0.8408 | 9.6488 | 15.2584 | 0.7051 | 13.5600 | 19.3862 | 0.4627 | |||
| T5Chem | 7.1283 | 11.2242 | 0.8385 | 6.7693 | 10.4041 | 0.8801 | 9.0982 | 14.3431 | 0.7665 | 13.4069 | 19.7521 | 0.6051 | |||
@@
| MFF | 6.4411 0.8157 | 9.6382 1.0206 | 0.8739 0.0292 |
| DRFP | 4.0995 0.1191 | 6.2424 0.2636 | 0.9474 0.0050 |
| Chemprop | 4.6430 0.1405 | 6.4306 0.1938 | 0.9441 0.0042 |
| YieldBert | 3.5532 0.1600 | 5.4480 0.3240 | 0.9599 0.0050 |
| T5Chem | 3.5059 0.1562 | 5.3181 0.2482 | 0.9662 0.0034 |
@@
5.2 Reaction Yield Prediction
Dataset. We use the Buchwald-Hartwig data set (Ahneman et al., 2018) to evaluate the performance of our model on the reaction yield prediction task. The dataset provides 10 random splits and four ligand-based out-of-sample splits for evaluation. The test sets under the four out-of-sample data splits contain reaction additives that are not included in the training sets. We use the raw data and the data split provided by Probst et al. and add the atom mapping for reactions according to the reaction template. In this study, we have standardized the yields in a range of 0 to 100. The statistical information of different splits is summarized in Appendix B.
Baselines. We compare our model with four powerful baselines: DRFP (Probst et al., 2022), Chemprop (Heid et al., 2024), YieldBert (Schwaller et al., 2021c) and T5Chem (Lu & Zhang, 2022) in the yield prediction task. The details of the baselines are presented in Appendix. E.
Implementation Details. It should be noted that the number of distinct reagents in the reaction conditions of the Buchwald-Hartwig dataset is less than 50, which makes it challenging to train a reaction condition encoder from scratch. Thus, we use a lightweight pretrained molecular encoder (Hu et al., 2020b) as our reaction condition encoder. For detailed information, please refer to Appendix C.4.
@@ 5.2 Reaction Yield Prediction
| MFF | 2.2923 0.0400 | 2.9388 0.0393 | 0.5891 0.0095 |
| DRFP | 0.9435 0.0293 | 1.2943 0.0387 | 0.9203 0.0042 |
| Chemprop | 0.3593 0.0099 | 0.5396 0.0299 | 0.9861 0.0015 |
| RXNFP | 0.3744 0.0094 | 0.5378 0.0204 | 0.9862 0.0010 |
| T5Chem | 0.6272 0.0184 | 0.8213 0.0226 | 0.9822 0.0011 |
@@
| MFF | 0.1421 0.0093 | 0.2122 0.0169 | 0.9055 0.0119 |
| DRFP | 0.1481 0.0094 | 0.2120 0.0147 | 0.9056 0.0117 |
| Chemprop | 0.1626 0.0118 | 0.2283 0.0174 | 0.8907 0.0130 |
| RXNFP | 0.1650 0.0111 | 0.2313 0.0197 | 0.8872 0.0193 |
| T5Chem | 0.1662 0.0108 | 0.2417 0.0182 | 0.8920 0.0121 |
@@
| Full Version | 26.91 | 39.80 | 44.82 | 50.73 | 3.6331 0.1259 | 5.5649 0.2839 | 0.9581 0.0049 | |
| - Atom Aligned Encoder | 25.69 | 39.00 | 44.51 | 50.77 | 3.7289 0.1291 | 5.7318 0.2520 | 0.9556 0.0044 | |
| - Reaction-Center-Aware Decoders | 26.64 | 39.24 | 44.14 | 50.33 | 3.6967 0.1710 | 5.7845 0.3948 | 0.9547 0.0066 | |
Performance Evaluation. The average results of ten random splits of the Buchwald-Hartwig dataset are summarized in Table 5 and the results of four out-of-sample splits are summarized in Table 4. Table 5 indicates that our model has achieved an average of 0.958 and an average MAE of 3.633 under ten random splits, placing it in the third position among all the compared methods. However, it is important to note that our model, which has not been pretrained on large-scale reaction data, is not significantly outperformed by two deep learning models that have. The difference in between our model and the strongest baseline is less than 0.01, and the gap in terms of MAE is less than 0.1. In the out-of-sample settings, our model demonstrates superior performance under various evaluation metrics. For instance, it attains an score of 0.898, surpassing the strongest baseline by 0.060. Notably, our model outperforms other baselines in all metrics, excluding the metric in the Test4 data split, by a significant margin. Furthermore, our model records an approximate 2.000 improvement in RMSE across all data splits.
@@
5.3 Reaction Selectivity Prediction
Datasets. We use the C-H functionalization dataset (Li et al., 2020) to demonstrate the predictive performance of regio-selectivity and use the experimental dataset (Zahrt et al., 2019; Li et al., 2023) regarding chiral phosphoric acid–catalyzed thiol addition to N-acylimines to illustrate the predictive performance of enantio-selectivity. There are 6114 chemical reactions in the C-H functionalization dataset, and the thiol addition dataset contains 43 catalysts and combinations of 5 5 reactants, which form 1075 reactions. These datasets were randomly divided into a training set and a test set with a 7: 3 ratio for 10 times.
Baseline. Considering that the reaction selectivity prediction task is fundamentally a regression problem, we have adapted three deep learning models originally designed for reaction yield prediction, Chemprop (Heid et al., 2024), T5chem (Lu & Zhang, 2022) and RXNFP (Schwaller et al., 2021b), to this task. In addition, two fingerprint-based methods, DRFP (Probst et al., 2022) and MFF (Sandfort et al., 2020), are compared. For the details of the baselines, please refer to Appendix E.
Performance Evaluation. The results for two datasets are summarized in Table 6 and Table 7. On the C-H functionalization dataset, our model achieves the best performance among all methods compared on all metrics, with an average MAE of 0.327, an average RMSE of 0.511, and an average of 0.988. Notably, our model outperforms two baselines RXNFP and T5chem, which are pretrained on large-scale reaction datasets. The results demonstrate that our model exhibits architectural superiority and a stronger understanding of chemical reactions compared to existing methods.